OpenStudy (anonymous):

Solve. (-7/3)(-2/21)(-13/14)(-3/13).

OpenStudy (akashdeepdeb):

Just multiply them simply. :)

OpenStudy (anonymous):

It told me the answer is wrong. I got -546/11466. I know I have to make the fraction smaller but I don't even know where to begin with that.

OpenStudy (akashdeepdeb):

Haha...No. Do not multiply everything like that. Cancel out some terms too! :D

OpenStudy (anonymous):

How? Can you please explain.

OpenStudy (akashdeepdeb):

I'll give you a hint. (216/343) * (49/36) Answer?

OpenStudy (akashdeepdeb):

When you see some terms divisible by some other just simply cancel them out \[\frac{216}{343} * \frac{49}{36}\] Now 216 is divisible by 216 And 343 is divisible by 49 Answer: 6/7

OpenStudy (akashdeepdeb):

Now try doing your question. :)

OpenStudy (anonymous):

How did you get 216/343 * 49/36?

OpenStudy (akashdeepdeb):

It is an EXAMPLE

OpenStudy (anonymous):

\(-\dfrac{7}{3}\times -\dfrac{2}{21} \times -\dfrac{13}{14} \times -\dfrac{3}{13}\) Start by finding a number in the numerator that is the same as a number in the denominator. Cancel that one out.

OpenStudy (anonymous):

@bobbj Your answer is correct. Try dividing 11466 by 546.